Sol Summary – December 27th

Sol 9

Ninth Day on Mars

David Murray

Mission Status: Everything is going well

Sol Activity Summary: The crew checked the generator, ate breakfast, and prepared for EVA. The EVA crew returned to the airlock to address complications with the radios. Once that was taken care of we completed the next half of the exercise. While we were finishing our EVA, we noticed a scratch in the helmet. We finished our task before returning to the airlock. After EVA, we completed our list of group tasks, wrote reports, and had dinner.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: Scratch noticed on helmet during EVA

Weather: Sunny

Crew Physical Status: Healthy

EVA: Navigation

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary, EVA report

Support Requested: None

Greenhab Report – December 27th

Greenhab Report

David Murray

27 Dec, 2017

Environmental control:

Ambient

Ambient with window/door open from 14:05 – 16:20

Working Hour: 14:14

Inside temp at working hour: 35 C

Outside temp during working hours: 7 C

Inside temperature high: 40 C

Inside temperature low: 19 C

Inside humidity: 46%

Heating

Functioning nominally

Turns on at 16 C and off at 20 C

Cooling

Not functional

Shade Cloth On/Off

On

The greenhab lights are on from 17:00 to 24:00 every night

Average temperature:

N/A

Changes to Crops

Cucumbers are wilting because of the heat.

Morning Research Observations

N/A

Daily Water Usage for Crops

6 gallons

Time(s) of Watering Crops

14:14

Aquaponics

N/A

Narrative

The cucumbers were wilting because the inside temperature got too hot and the soil was dry. They recieved extra water today and they should be fine by tomorrow.

Support/Supplies Needed

N/A

 

Operations Report – December 27th

SOL: 9

Name of person filing report: A.Passeron

Non-nominal systems:

– Helmet :

At 12:17pm, one of our crew accidentally tripped over a rock and fell forward during the EVA, scratching his helmet and our camera. This is not a crack. No medical injury, as determined by HSO.

Generator (hours run):

Generator turned off at 8:00am

Start charging battery at 8:00am

Battery charge is 80% at 5:25pm

Generator turned on at 5:25pm

Solar— SOC

@ 8:00am : 100%

@ 9:40am : 89%

@ 2:50pm : 100%

@ 5:25pm : 85%

@ 7:00pm : 85%

Diesel – 76%

Propane – 46%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2.5 Gallons (1 container remaining)

Water (trailer) – 20 gallons

Water (static) – 250 gallons

Trailer to Static Pump used – No

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: No

ATVs Used: No

Oil Added?

ATV Fuel Used:

# Hours the ATVs were used today: 00:00 hours

Notes on ATVs: ATVs were nominal.

Deimos rover used: No

Number of hours: 95.2

Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Number of hours: 4.4

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Number of hours: 5.2

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Number of hours: 4.3

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: No

Number of hours: 1.7

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: One of our crew accidentally tripped over a rock and fell forward during today’s EVA, scratching his helmet. This is not a crack.

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: